The following information is from Harrelson Inc.'s, financial statements. Sales (all credit) were $21.60 million for last year.    

Sales to total assets 1.80 times
Total debt to total assets 40 %
Current ratio 3.00 times
Inventory turnover 4 times
Average collection period 25 days
Fixed asset turnover 5 times

Complete the balance sheet: (Use a 360-day year. Do not round intermediate calculations. Input your answers in millions rounded to 2 decimal places.)

$ in millions $ in millions
Cash Current debt
Accounts receivable Long-term debt
Inventory Total debt
Total current assets
Fixed assets Equity
Total assets Total debt and stockholders' equity
Answer #1
Ans. ASSETS in millions LIABILITIES in millions
Cash 0.78 Current debt 2.56
Accounts receivables 1.5 Long - term debt 2.24
Inventory 5.4 Total debt 4.8
Total current assets 7.68
Fixed assets 4.32 Equity 7.2
Total assets 12 Total debt and stockholders' equity 12
*Calculations:
1 Sales to Total assets =   Sales / Total assets
1.80   = 21.60 / Total assets
Total assets = 21.60 / 1.80
$12 million
2 Total debt to Total assets = Total debt / Total assets
0.40   =    Total debt / 12
Total debt   =   12 * 0.40
$4.8 million
3 Fixed assets turnover =   Sales / Fixed assets
5   = 21.60 / Fixed assets
Fixed assets = 21.60 / 5
$4.32 million
4 Current assets = Total assets - Fixed assets
12 - 4.32
$7.68 million
5 Current ratio = Current assets / Current liabilities
3 = 7.68 / Current liabilities
Current liabilities = 7.68 / 3
$2.56 million
6 Total debt = Current liabilities + Long term debt
4.8 = 2.56 + Long term debt
Long term debt = 4.8 - 2.56
$2.24 million
7 Inventory turnover = Sales / Inventory
4 = 21.60 / inventory
Inventory = 21.60 / 4
$5.4 million
8 Average collection period = No. of days in year / Receivables turnover
25 =   360 / Receivables turnover
Receivables turnover = 360 / 25
14.4 times
Receivables turnover = Sales / Receivables
14.4 = 21.60 / Receivables
Receivables   = 21.60 / 14.4
$1.5 million
9 Current assets = Inventory + Receivables + Cash
7.68 = 5.4 + 1.5 + Cash
Cash = 7.68 - 5.4 - 1.5
$0.78 million
10 Total debt and stockholder's equity = Total debt + Equity
12 = 4.80 + Equity
Equity = 12 - 4.80
Equity   = 7.20
*Total debt and stockholder's equity = Total assets